Circular No. 7603

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ag IN MCG +08-18-9 T. Puckett and M. Peoples, Mountain Town, GA, report the discovery

of an apparent supernova (mag 18.2) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Mar. 27.27 UT. SN 2001ag is located at α = 9^{h}34^{m}31^{s}.01, δ = +46^{o}27′38′′.7 (equinox 2000.0), which is 5′′.9 west and 3′′.0 south of the center of MCG +08-18-9. The new object was also present on an unfiltered CCD frame taken on Mar. 28.03, but it was not present on a frame taken on Jan. 3.42, neither does it appear on Palomar Sky Survey images taken on 1990 Jan. 24, 1995 Nov. 23 (limiting mag ∼ 21.0), or 1953 Feb. 14 (limiting mag ∼ 20.0).

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001ag, obtained by P. Berlind on Mar. 28.19 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. The spectrum consists of a blue continuum with well-developed P-Cyg lines of hydrogen. Narrow Hα emission superposed on the spectrum of the supernova indicates a recession velocity of 8000 km/s for the host galaxy. The expansion velocity derived from the minimum of the Hβ line is ∼ 9100 km/s.

SUPERNOVA 2001ah IN UGC 6211 Puckett and Peoples also report the discovery of another apparent

supernova (mag 18.3) on an unfiltered CCD frame (limiting mag 20.5) taken as above on Mar. 27.24 UT. SN 2001ah is located at α = 11^{h}10^{m}29^{s}.82, δ = +55^{o}09′39′′.0 (equinox 2000.0), which is 4′′.3 west and 32′′.4 south of the center of UGC 6211. The new object was also present on an unfiltered CCD frame taken on Mar. 28.11. SN 2001ah does not appear on Palomar Sky Survey images taken on 1991 Apr. 4, 1991 Feb. 10 (limiting mag ∼ 21.0), or 1953 Feb. 14 (limiting mag ∼ 20.0). M. Schwartz, Cottage Grove, OR; and W. Li, University of California at Berkeley, report the independent discovery by LOTOSS (cf. IAUC 7514) of this object at mag ∼ 17.8 on unfiltered images taken with the 0.5-m Tenagra III automatic telescope on Mar. 27.2 and 28.3. They give position end figures 29^{s}.85, 39′′.0; an image taken on Mar. 9.2 showed nothing at this position (limiting mag 19.0).

2001 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green